Precious, Inc. is a merchandiser of a single line of golden rings. At the beginning of the day, the shop had 10 rings in its inventory. During the day, 4 new rings were delivered to the shop. By close of business, only 8rings remained in inventory. The purchase price of each ring from the supplier is $276. In addition, the company pays $5 for shipping and delivery insurance on each ring that they purchase. What is the companys Gross Profit for the day if it sells each ring for $607? The Gross Profit for the day is$
Answer #1

Number of units sold = Beginning inventory + Purchases - Ending inventory

= 10 + 4 - 8

= 6

Sales = Number of units sold x Selling price per unit

= 6 x 607

= $3,642

Purchase price per ring = $276

Shipping and delivery insurance = $5 per ring

Hence, total cost per ring = 276 + 5

= $281

Cost of goods sold = Number of units sold x Cost per ring

= 6 x 281

= $1,686

Gross profit = Sales - Cost of goods sold

= 3,642 - 1,686

= $1,956

Kindly give a positive rating if you are satisfied with the answer. Feel free to ask if you have any doubts. Thanks